# Decoupled Global-Local Alignment for Improving Compositional Understanding

### v1

@@ Abstract.
Contrastive Language-Image Pre-training (CLIP) has achieved success on multiple downstream tasks by aligning image and text modalities. However, the nature of global contrastive learning limits CLIP’s ability to comprehend compositional concepts, such as relations and attributes. Although recent studies employ global hard negative samples to improve compositional understanding, these methods significantly compromise the model’s inherent general capabilities by forcibly distancing textual negative samples from images in the embedding space. To overcome this limitation, we introduce a Decoupled Global-Local Alignment (DeGLA) framework that improves compositional understanding while substantially mitigating losses in general capabilities. To optimize the retention of the model’s inherent capabilities, we incorporate a self-distillation mechanism within the global alignment process, aligning the learnable image-text encoder with a frozen teacher model derived from an exponential moving average. Under the constraint of self-distillation, it effectively mitigates the catastrophic forgetting of pretrained knowledge during fine-tuning. To improve compositional understanding, we first leverage the in-context learning capability of Large Language Models (LLMs) to construct about 2M high-quality negative captions across five types. Subsequently, we propose the Image-Grounded Contrast (IGC) loss and Text-Grounded Contrast (TGC) loss to enhance vision-language compositionally. Extensive experimental results demonstrate the effectiveness of the DeGLA framework. Compared to previous state-of-the-art methods, DeGLA achieves an average enhancement of 3.5% across the VALSE, SugarCrepe, and ARO benchmarks. Concurrently, it obtains an average performance improvement of 13.0% on zero-shot classification tasks across eleven datasets. Our code will be released at https://github.com/xiaoxing2001/DeGLA.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; April; 2025isbn: 978-1-4503-XXXX-X/YY/MMccs: Computing methodologies Computer vision representations
@@ 1. Introduction
The rapid expansion of mobile networks and social platforms significantly boosts the large-scale generation of image-text pairs, laying a crucial foundation for vision-language pre-training. In particular, Contrastive Language-Image Pre-training (Radford et al., 2021a) (CLIP) leverages two distinct unimodal encoders for images and texts, and employs contrastive loss (Oord et al., 2018a) for representation learning. After pretraining on extensive image-text pairs sourced from the internet, CLIP exhibits strong transferability and is extensively applied across various tasks, including image captioning (Alayrac et al., 2022; Li et al., 2023b; Mañas et al., 2022), object detection (Kaul et al., 2023; Ma et al., 2023; Wu et al., 2024b), and segmentation (Li et al., 2022b; Cho et al., 2024; Yu et al., 2023; Minderer et al., 2022; Hoyer et al., 2025; Gu et al., 2021; Zang et al., 2022; Wu et al., 2024a).
@@ 2.1. Vision-Language Contrastive Learning
2.2. Vision-Language Compositionality
@@ 3. Method
In subsequent sections, we first present the preliminary knowledge of CLIP in Section 3.1, followed by a detailed exposition of our proposed LLM-driven negative caption generation pipeline in Section 3.2. A comprehensive description of the DeGLA training framework is provided in Section 3.3.
@@ 4.4. Analysis
Ablation of Losses. We analyze the impact of the weights $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ of the three losses on the performance of ARO, as illustrated in Figure 6(a) and Figure 6(b). Initially, we identify the optimal combination of ($\lambda_{1},\lambda_{2}$), finding that the model achieves its best performance, 86.1%, with $\lambda_{1}$ = 0.1 and $\lambda_{2}$ = 0.1. Consequently, we adopt these parameters. Subsequent analysis of $\lambda_{3}$ variations reveals that increasing distillation enhances zero-shot classification performance but reduces combinatorial understanding, reflecting the conflicting objectives of self-distillation and compositional reasoning. Ultimately, we select $\lambda_{3}$ = 0.005, as this setting provides a well-balanced model, as indicated by the red dashed circle in the figure.
@@ 4.5. Case Study
We provide a case study comparison between CLIP and DeGLA on ARO and SugarCrepe in Figure 7. The results show that the CLIP model exhibits a pronounced “bag-of-words” phenomenon, as indicated in previous work NegCLIP (Yuksekgonul et al., 2022), meaning it struggles with recognizing structural changes in texts, such as swapping the order of targets in a sentence, adding targets, or replacing targets. For instance, CLIP fails to select the correct text from options like “Two ducks swim in a pond with green water.” and “Two swans swim in a pond with green ducks.”. In contrast, DeGLA is capable of correctly distinguishing between them based on the given image. Notably, after training on our diverse compositionality-related hard negative data generated by our proposed pipeline in Section 3.2, DeGLA demonstrates enhanced compositional reasoning understanding compared to pretrained CLIP, such as intra-sentential reshuffle and minimal semantic substitution, thereby underscoring the effectiveness of our data generation pipeline.
@@ 5. Conclusion
In this paper, we observe that while previous methods enhance CLIP’s compositional understanding, they often compromise its general capabilities. To overcome this limitation, we introduce a Decoupled Global-Local Alignment (DeGLA) framework, which not only improves compositional understanding but also significantly reduces losses in general capabilities. To optimize the retention of the model’s inherent capabilities, we incorporate a self-distillation mechanism within the global alignment process, aligning the learnable image-text encoder with a frozen teacher model derived from an exponential moving average. For enhancing compositional understanding, we leverage the in-context learning capabilities of Large Language Models (LLMs) to generate approximately 2M high-quality negative captions across five types. We further propose the Image-Grounded Contrast (IGC) loss and Text-Grounded Contrast (TGC) loss to improve vision-language compositionality. Extensive experimental results demonstrate the effectiveness of the DeGLA framework. We hope that our work provides insights into vision-language compositional understanding.
@@ A.1. Training Details
Training GPUs 8×8$\times$8 ×V100

### v2

@@ Abstract.
Contrastive Language-Image Pre-training (CLIP) has achieved success on multiple downstream tasks by aligning image and text modalities. However, the nature of global contrastive learning limits CLIP’s ability to comprehend compositional concepts, such as relations and attributes. Although recent studies employ global hard negative samples to improve compositional understanding, these methods significantly compromise the model’s inherent general capabilities by forcibly distancing textual negative samples from images in the embedding space. To overcome this limitation, we introduce a Decoupled Global-Local Alignment (DeGLA) framework that improves compositional understanding while substantially mitigating losses in general capabilities. To optimize the retention of the model’s inherent capabilities, we incorporate a self-distillation mechanism within the global alignment process, aligning the learnable image-text encoder with a frozen teacher model derived from an exponential moving average. Under the constraint of self-distillation, it effectively mitigates the catastrophic forgetting of pretrained knowledge during fine-tuning. To improve compositional understanding, we first leverage the in-context learning capability of Large Language Models (LLMs) to construct about 2M high-quality negative captions across five types. Subsequently, we propose the Image-Grounded Contrast (IGC) loss and Text-Grounded Contrast (TGC) loss to enhance vision-language compositionally. Experimental results across both general and compositional reasoning tasks validate the effectiveness of the DeGLA framework. Our code is released at https://github.com/xiaoxing2001/DeGLA.
@@
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755032isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer vision representations
@@ 1. Introduction
The rapid expansion of mobile networks and social platforms significantly boosts the large-scale generation of image-text pairs, laying a crucial foundation for vision-language pre-training (Gu et al., 2025b, a). In particular, Contrastive Language-Image Pre-training (Radford et al., 2021a) (CLIP) leverages two distinct unimodal encoders for images and texts, and employs contrastive loss (Oord et al., 2018a) for representation learning. After pretraining on extensive image-text pairs sourced from the internet, CLIP exhibits strong transferability and is extensively applied across various tasks, including image captioning (Alayrac et al., 2022; Li et al., 2023b; Mañas et al., 2022), object detection (Kaul et al., 2023; Ma et al., 2023; Wu et al., 2024b), and segmentation (Li et al., 2022b; Cho et al., 2024; Yu et al., 2023; Minderer et al., 2022; Hoyer et al., 2025; Gu et al., 2021; Zang et al., 2022; Wu et al., 2024a).
@@ 2.1. Vision-Language Contrastive Learning
$2.$ Vision-Language Compositionality
@@ 3. Method
In subsequent sections, we first present the preliminary knowledge of CLIP in Section 3.1, followed by a detailed exposition of our proposed LLM-driven negative caption generation pipeline in Section 3.2. A comprehensive description of the DeGLA training framework is provided in Section $3.$
@@ 4.5. Case Study
Figure 7 compares CLIP and DeGLA on ARO and SugarCrepe. Consistent with NegCLIP (Yuksekgonul et al., 2022), CLIP shows a “bag-of-words” issue, struggling with structural text changes like word order swaps or substitutions—for example, failing to distinguish between “Two ducks swim in a pond with green water” and “Two swans swim in a pond with green ducks.” In contrast, DeGLA correctly differentiates these based on the image. After training on our compositionality-focused hard negatives (Section 3.2), DeGLA demonstrates improved compositional reasoning over pretrained CLIP, validating the effectiveness of our data generation pipeline.
@@ 5. Conclusion
In this paper, we observe that while previous methods enhance CLIP’s compositional understanding, they often compromise its general capabilities. To overcome this limitation, we introduce a Decoupled Global-Local Alignment (DeGLA) framework, which not only improves compositional understanding but also significantly reduces losses in general capabilities. To optimize the retention of the model’s inherent capabilities, we incorporate a self-distillation mechanism within the global alignment process, aligning the learnable image-text encoder with a frozen teacher model derived from an exponential moving average. For enhancing compositional understanding, we leverage the in-context learning capabilities of Large Language Models (LLMs) to generate approximately 2M high-quality negative captions across five types. We further propose the Image-Grounded Contrast (IGC) loss and Text-Grounded Contrast (TGC) loss to improve vision-language compositionality.
Acknowledgements.
This work is supported by National Natural Science Foundation of China (Grant No. 62301046)
@@ A.1. Training Details
Training GPUs $8\timesV$100